2. In the scenario in assignment 1, the deputy county manager wants the cost allocation to be redone using expenditures as the base. What are the new cost allocations?

Expenditures FTE's
Deputy Co. Manager $250,000 2
Human Resources $800,000 9
Information Technology $3,525,000 32
Transportation $37,525,813 275
Facilities $28,451,254 150
Solid Waste $8,752,486 24
Stormwater Control $35,248,125 412

The answer's from question one were Allocation = 114552678 / 904 *2 = 253435.13 for deputy county manager

1,140,458.08 for human resources and so on.
Answer #1

First we calculate per FTE cost by dividing total expenditures by total FTE's.

Total expenditure = $250,000+$800,000+$3,525,000+$37,525,813+$28,451,254+$8,752,486+$35,248,125 = $114,552,678

Total FTE's = 2+9+32+275+150+24+412 = 904

per FTE expenditure = $114,552,678/904 = $126,717.564159292‬

New cost allocations are as below:

New cost allocation = per FTE expenditure*no. of FTE's

Deputy Co. Manager = $126,717.564159292‬*2 = $253,435.13

Human Resources = $126,717.564159292*9 = $1,140,458.08

Information Technology = $126,717.564159292*32 = $4,054,962.05

Transportation = $126,717.564159292*275 = $34,847,330.14

Facilities = $126,717.564159292*150 = $19,007,634.62

Solid Waste = $126,717.564159292*24 = $3,041,221.54

Storm water Control = $126,717.564159292*412 = $52,207,636.43